102. Time swinging. The period T (time in seconds for one complete cycle) of a simple pendulum is related to the length L (in feet) of the pendulum by the formula 8T^2 = pie ^2L. If a child is on a swing with 10 feet chain, then how long does it take to complete one cycle? my brain is dead cam some one help my brain rejuvenate :)
Answer by tutorcecilia(2152)

8T^2 = pie ^2L
8T^2 = pie ^2(10) [Substitute 10 for the length (L) and solve]
T^2=10pi^2/8 [solve for T]
T^2=12.337
sqrt(T^2)=sqrt(12.337)
T=3.5124 seconds
.
Check by plugging all of the values back into the original problem and solve]